Lottery Prize (millions) $                           14.00 millions
Annuity (millions) $                             1.00 per year
APR ( R ) 7.00% annual
N (begin now) in years                               25.00 years
Benefit of lumpsum option (Excel) Gain/Loss of ???? millions
Benefit of lumpsum option (VBA) Gain/Loss of ???? millions
Benefit of lumpsum (Math, Extra Credit) Gain/Loss of ???? millions

Juan has just won a lottery. He can get $14 million now (after-tax, lumpsum or one-time option), or he can get $1 million annually for 25 years starting today (after-tax, 25 payments of $1 million each). Whichever option he chooses, he can save his money at a local bank, which is offering a 7% per year return on deposits.
Juan is not sure if he should pick a lumpsum option or 25-payments option. He comes to you for help. What is the benefit (or loss) of choosing the lumpsum option in today’s dollars?
Answer #1

Present value of lump-sum option = $14 million

Present value of annuity = 1+1.07-1+1.07-2+1.07-3+.........+1.07-24= (1 - 1.07-25) / (1 - 1.07-1) =$12.469334 million

As the PV of lump- sum option is higher, Juan should opt for the lump-sum option with a benefit of

$(14 - 12.469334) million=$1530666 in today's dollars.